OpenStudy (anonymous):

carol works 5 days a week for 6 hours each day and earns 6.50 per hour. she saves 18% of her weekly paycheck for college expenses. if she takes a two week vacation during the summer, how much will she save

OpenStudy (wikiemol):

well how many weeks does she work?

OpenStudy (anonymous):

idkk lol thats what the question says

OpenStudy (anonymous):

my teacher got 1755 as the answer

OpenStudy (wikiemol):

well I suppose you can say p = (5*6*6.5)w and so her savings are p*0.18

OpenStudy (wikiemol):

I dont know how many weeks carol works though.

OpenStudy (anonymous):

yeah not clear right?

OpenStudy (anonymous):

maybe you are supposed to say there are 52 weeks in a year, so she works the other 50

OpenStudy (anonymous):

then her total pay is \(50\times 5\times 6\times 6.50\)

OpenStudy (anonymous):

fifty weeks, five days a week, six hours a day, 6.50 per hour

OpenStudy (anonymous):

my teacher got 1755 but i got 35.10

OpenStudy (anonymous):

girlfriend needs a new job

OpenStudy (anonymous):

i get \[50\times 5\times 6\times 6.50=9750\] and \[.18\times 7950=1755\]

OpenStudy (anonymous):

oo thanks ur soo smart lol!